I can't understand what does it mean!
Mrs Brown, who is 51 years old, has a daughter 21 years old. How long will it be before mother is twice as old as her daughter

Answer by Earlsdon(5028)

Let x = the number of years hence when Mrs. Brown (B, now 51) will be twice her daughter's age (D, now 21).
M = 51
D = 21
M+x = 2(D+x) Substitute M = 51 and D = 21.
51+x = 2(21+x)
51+x = 42+2x Subtract x from both sides.
51 = 42+x Subtract 42 from both sides.
9 = x
In 9 year's time, Mrs. Brown will be twice her daughter's age.
Check:
M+x = 2(D+x)
51+9 = 2(21+9)
60 = 2(30) OK!